# Convergence order of the quantization error for self-affine measures on Lalley-Gatzouras carpets

### v1

@@ Abstract.
convergence order, quantization error, self-affine measures, Lalley-Gatzouras carpets.
@@ 1. Introduction
The quantization problem consists in the approximation of a given probability measure by discrete probability measures of finite support in $L_{r}$-metrics. We refer to Graf and Luschgy [10] for rigorous mathematical foundations of quantization theory and [16] for its deep background in information theory and some engineering technology. Applications of quantization theory in numerical integration and mathematical finance have been investigated by Gruber [17], Pagès and his coauthors (cf. [8, 27]). A recent breakthrough by Kesseböhmer et al reveals the close connection between quantization theory and the $L^{q}$-spectra for compactly supported probability measures on $\mathbb{R}^{d}$ [21]. One may see [11, 12, 13, 14, 15, 20, 25, 29, 31, 32] for more results on the quantization for fractal measures.
@@ 1.1. Quantization error and its asymptotics
Let $|x|$ denote the Euclidean norm of $x\in\mathbb{R}^{d}$. Given $r\in(0,\infty)$, let $\nu$ be a Borel probability measure on $\mathbb{R}^{d}$ with $\int|x|^{r}d\mu(x)<\infty$. We denote by ${\rm card}(B)$ the cardinality of a set $B$. For every $n\geq 1$, we write $\mathcal{D}_{n}:=\{\alpha\subset\mathbb{R}^{d}:1\leq{\rm card}(\alpha)\leq n\}$. For $x\in\mathbb{R}^{d}$ and a set $A\subset\mathbb{R}^{d}$, let $d_{A}(x):=\inf_{y\in A}d(x,y)$. For a Borel function $g$, let $\|g\|_{r}:=(\int|g|^{r}d\nu)^{1/r}$. The $n$th quantization error for $\nu$ of order $r$ can be defined by
@@
By [10, Lemma 3.4], $e_{n,r}(\nu)$ is equal to the minimum error in the approximation of $\nu$ with discrete probability measures supported at most $n$ points, in the $L_{r}$-metric. We refer to [10] for various interpretations of $e_{n,r}(\nu)$ in different contexts.
@@
The upper (lower) quantization dimension $\overline{D}_{r}(\nu)\;(\underline{D}_{r}(\nu))$ for $\nu$ of order $r$, is (typically) the critical point at which the upper (lower) quantization coefficient jumps from infinity to zero (cf. [10, 29]). By [10, Proposition 11.3], we have
@@
When $\underline{D}_{r}(\nu)=\overline{D}_{r}(\nu)$, we say that the quantization dimension exists and denote the common value by $D_{r}(\nu)$.
@@
Assuming the open set condition, Graf and Luschgy established complete results for the asymptotics of the quantization error for self-similar measures on $\mathbb{R}^{d}$ (cf. [11]). Their methods and results have greatly influenced almost all the subsequent study on the quantization for fractal measures.
While it is not easy even to estimate the upper and lower quantization dimension for general probability measures, Kesseböhmer et al [21], identified the upper quantization dimension for an arbitrary compactly supported measure with its Rényi dimension at a critical point and provided several sufficient conditions that guarantee the existence of the quantization dimension. This work, along with Feng-Wang’s results in [7], yields that the quantization dimension exists for the self-affine measures on a large class of planar self-affine sets, including Lalley-Gatzouras carpets. Further, the quantization dimension can be implicitly expressed in terms of Feng-Wang’s formula [7, Theorem 2].
However, the results in [21] do not provide us with useful information on the asymptotic order for the quantization error. In the present paper, we will determine the exact convergence order of the quantization error for the self-affine measures on Lalley-Gatzouras carpets in general.
@@ 1.2. Lalley-Gatzouras carpets
Let $G:=\{(i,j):\;1\leq i\leq n_{j},1\leq j\leq m\}$ and $G_{y}:=\{1,\ldots,m\}$. We consider
(1.3) $f_{ij}(x,y)=(x,y)\left(\begin{array}[]{cccc}a_{ij}&0\\ 0&b_{j}\\ \end{array}\right)+(c_{ij},d_{j}),\;(x,y)\in\mathbb{R}^{2};\;(i,j)\in G.$
With the assumptions (A1)-(A4), the iterated function system $\{f_{ij}\}_{(i,j)\in G}$ satisfies the open set condition with respect to $U:=(0,1)^{2}$: $f_{ij}(U),(i,j)\in G$, are disjoint subsets of $U$. By [18], there exists a unique non-empty compact set $E$ satisfying
@@
We call $\mu$ the self-affine measure associated with $\mathcal{I}$ and $\mathcal{P}$. Let $E_{0}:=[0,1]^{2}$. For $l\geq 1$ and $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\in G^{l}$, we write $f_{\sigma}:=f_{i_{1}j_{1}}\circ\cdots\circ f_{i_{l}j_{l}}$. We will call $f_{\sigma}(E_{0})$ a cylinder of order $l$.
In the past decades, self-affine sets and self-affine measures have attracted great interest of mathematicians (cf. [1, 2, 4, 6, 7, 19, 23, 20, 24, 26, 28]). As was noted in the literature, problems concerning typical self-affine sets and self-affine measures are usually difficult. In [7], D.-J. Feng and Y. Wang determined various dimensions for a class of planar self-affine sets and self-affine measures, and established several computable formulas for such measures. These results were then generalized to more general self-affine measures on the plain by Fraser [9] and were further generalized by Kolossváry [22] to self-affine measures on sponges in $\mathbb{R}^{d}$ by introducing a novel pressure function.
@@
$G\subset\{0,1,\ldots,n_{0}-1\}\times\{0,1,\ldots,m_{0}-1\};$
$a_{ij}\equiv\frac{1}{n_{0}}\;((i,j)\in G),\;b_{j}\equiv\frac{1}{m_{0}}\;(j\in G_{y}),$
$c_{ij}\equiv\frac{i}{n_{0}}\;((i,j)\in G),\;d_{j}\equiv\frac{j}{m_{0}}\;(j\in G_{y});$
where $G_{y}$ denotes the projection of $G$ onto the $y$-axis. Then the Lalley-Gatzouras carpet $E$ degenerates to a Bedford-McMullen carpet. Let $\xi=[\frac{\log m_{0}}{\log n_{0}}]$ and $q_{j}:=\sum_{i:(i,j)\in G}p_{ij}$. Kesseböhmer and Zhu proved that the quantization dimension $d_{r}$ for $\mu$ of order $r$ is implicitly given by (cf. [20])
$\bigg{(}\sum_{(i,j)\in G}(p_{ij}m_{0}^{-r})^{\frac{d_{r}}{d_{r}+r}}\bigg{)}^{\xi}\bigg{(}\sum_{j\in G_{y}}(q_{j}m_{0}^{-r})^{\frac{d_{r}}{d_{r}+r}}\bigg{)}^{1-\xi}=1.$
@@
Thanks to the relatively fine structure of Bedford-McMullen carpets, we considered the auxiliary coding space $\Phi_{\infty}:=G^{\mathbb{N}}\times G_{y}^{\mathbb{N}}$ and constructed a Bernoulli product $W$ as an auxiliary measure. Then the upper and lower quantization coefficient for $\mu$ were well estimated via the measure $W$, by going back and forth between $E$ and $\Phi_{\infty}$. In the Lalley-Gatzouras case, however, neither the linear parts nor the translations of the mappings $f_{ij},(i,j)\in G$, need to be constant. This substantially adds to the complexity of the structure of the carpets and the difficulty in analyzing the quantization error.
Combining [7, Theorem 2] and [20, Theorem 1.11], we obtain that, the quantization dimension exists for an arbitrary self-affine measure on Lalley-Gatzouras carpets; and the quantization dimension can be expressed in terms of the $L^{q}$-spectrum of the projection of $\mu$ onto the $y$-axis. However, such expressions seem to be inconvenient in study of the asymptotic order of the quantization error for $\mu$. We will establish a new expression for the quantization dimension in terms of cylinders of the same order. This expression is more convenient for us to construct suitable auxiliary measures which will be used to estimate the upper and lower quantization coefficient.
@@ 1.3. Some notations
For $\sigma=((i_{1},j_{1}),\ldots,(i_{k},j_{k}))\in G^{k}$ and $1\leq h\leq k$, we define $|\sigma|=k$ and $\sigma|_{h}:=((i_{1},j_{1}),\ldots,(i_{h},j_{h}))$. Let $k\geq 1$. Let $\theta$ denote the empty word. For $\sigma\in G^{1}$, we define $\sigma^{-}=\theta$. For every $k\geq 2$ and $\sigma\in G^{k}$, we define $\sigma^{-}:=\sigma|_{k-1}$. Let $G^{*}:=\bigcup_{k\geq 1}G^{k}$. If $\sigma,\omega\in G^{*}\cup G^{\mathbb{N}}$ satisfy $\sigma=\omega|_{|\sigma|}$, we say that $\sigma$ is comparable with $\omega$ and write $\sigma\preceq\omega$. If $\sigma\preceq\omega$ and $\sigma\neq\omega$, we write $\sigma\precneqq\omega$. We call $\sigma,\omega\in G^{*}$ incomparable if neither $\sigma\preceq\omega$ nor $\omega\preceq\sigma$. For the words $\tau$ in $G_{y}^{*}:=\bigcup_{k\geq 1}G_{y}^{k}$, we define $\tau|_{h},\tau^{-}$ and the partial order $\preceq$, in a similar manner.
@@
$\omega_{y}:=(j_{1},\ldots,j_{l}),\;(\sigma\ast\tau)_{y}=(j_{1},\ldots,j_{k});$
$a_{\omega}:=\prod_{h=1}^{l}a_{i_{h}j_{h}},\;b_{\tau}=\prod_{h=1}^{k}b_{j_{h}};\;a_{\theta}=b_{\theta}:=1.$
@@
$B_{L,\sigma}:=d_{j_{1}}+\sum_{p=2}^{k}\bigg{(}\prod_{h=1}^{p-1}b_{j_{h}}\bigg{)}d_{j_{p}},\;A_{R,\sigma}:=B_{L,\sigma}+\prod_{h=1}^{k}b_{j_{h}}.$
@@
For $\sigma,\omega\in\Psi^{*}$, we write $\sigma\preceq\omega$ if $F_{\sigma}\supset F_{\omega}$. In typical Lalley Gatzouras case, the diameters of approximate squares of the same order can be widely different.
We will also consider the auxiliary coding space $\Phi_{\infty}=G^{\mathbb{N}}\times G_{y}^{\mathbb{N}}$. We define
@@
We write $\mathcal{L}(\sigma)_{L}:=\sigma_{L}$ and $\mathcal{L}(\sigma)_{R}:=\sigma_{R}$. Since confusion could arise, sometimes we simply write $\sigma$ for $\mathcal{L}(\sigma)$. For $\sigma,\omega\in\Phi^{*}$, we write $\sigma\preceq\omega$, if $\sigma_{L}\preceq\omega_{L}$ and $\sigma_{R}\preceq\omega_{R}$. We call $\sigma,\omega\in\Phi^{*}$ incomparable if neither $\sigma\preceq\omega$ nor $\omega\preceq\sigma$. The difference between $\Psi^{*}$ and $\Phi^{*}$ lies in the fact that they are endowed with different partial orders. For $\sigma\in\Phi^{*}$, we define $\sigma]:=[\sigma_{L}]\times[\sigma_{R}]$, where
@@
Let $|B|$ denote the diameter of a set $B\subset\mathbb{R}^{2}$ and $B^{\circ}$ its interior.
@@
    For $\sigma,\omega\in\Psi^{*}$, $F_{\sigma},F_{\omega}$ are either non-overlapping (i.e. $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset)$, or one of them is a subset of the other.

### v10

@@ Abstract.
convergence order, quantization error, $L^{q}$-spectrum, self-affine measures, Lalley-Gatzouras carpets.
@@ 1. Introduction
The quantization problem consists in the approximation of a given probability measure by discrete probability measures of finite support in $L_{r}$-metrics. We refer to Graf and Luschgy [5] for rigorous mathematical foundations of quantization theory and [10] for its deep background in information theory and some engineering technology. One may also see [6, 7, 8, 9, 13, 18, 24, 25, 27, 28] for more results on the quantization for fractal measures.
@@ 1.1. Quantization error and its asymptotics
Let $|x|$ denote the Euclidean norm of $x\in\mathbb{R}^{d}$. Given $r\in(0,\infty)$, let $\nu$ be a Borel probability measure on $\mathbb{R}^{d}$ with $\int|x|^{r}d\mu(x)<\infty$. We denote by ${\rm card}(B)$ the cardinality of a set $B$. For every $n\in\mathbb{N}$, we write $\mathcal{D}_{n}:=\{\alpha\subset\mathbb{R}^{d}:1\leq{\rm card}(\alpha)\leq n\}$. For $x\in\mathbb{R}^{d}$ and a set $A\subset\mathbb{R}^{d}$, let $d_{A}(x):=\inf_{y\in A}d(x,y)$. For a Borel measurable function $g$ on $\mathbb{R}^{d}$, let $\|g\|_{r}:=(\int|g|^{r}d\nu)^{\frac{1}{r}}$. The $n$th quantization error for $\nu$ of order $r$ can be defined by
@@
By [5, Lemma 3.4], $e_{n,r}(\nu)$ is equal to the minimum error in the approximation of $\nu$ with discrete probability measures supported at most $n$ points, in the $L_{r}$-metric. We refer to [5] for various interpretations of $e_{n,r}(\nu)$ in different contexts.
@@
The upper (lower) quantization dimension $\overline{D}_{r}(\nu)\;(\underline{D}_{r}(\nu))$ for $\nu$ of order $r$, is (typically) the critical point at which the upper (lower) quantization coefficient jumps from infinity to zero (cf. [5, 24]). By [5], we have
@@
If $\underline{D}_{r}(\nu)=\overline{D}_{r}(\nu)$, we say that the quantization dimension exists and denote the common value by $D_{r}(\nu)$.
@@
1.2. Some known results on the quantization problem
Let $(f_{i})_{i=1}^{N}$ be a set of contractive similarities on $\mathbb{R}^{d}$ with contraction ratios $(c_{i})_{i=1}^{N}$. We say that $(f_{i})_{i=1}^{N}$ satisfies the open set condition (OSC), if there exists a bounded non-empty open set $U$ such that $f_{i}(U),1\leq i\leq N$, are pairwise disjoint subsets of $U$. According to [11], there exists a unique non-empty compact set $F$ satisfying $F=\bigcup_{i=1}^{N}f_{i}(F)$. We call $F$ the self-similar set determined by $(f_{i})_{i=1}^{N}$. Given a positive probability vector $(p_{i})_{i=1}^{N}$, there exists a unique Borel probability measure satisfying $\nu=\sum_{i=1}^{N}p_{i}\nu\circ f_{i}^{-1}$. This measure is called the self-similar measure associated with $(f_{i})_{i=1}^{N}$ and $(p_{i})_{i=1}^{N}$. Let $\xi_{r}$ be implicitly defined by $\sum_{i=1}^{N}(p_{i}c_{i}^{r})^{\frac{\xi_{r}}{\xi_{r}+r}}=1$. Assuming the OSC for $(f_{i})_{i=1}^{N}$, Graf and Luschgy (cf. [6]) proved that
$0<\underline{Q}^{\xi_{r}}_{r}(\nu)\leq\overline{Q}^{\xi_{r}}_{r}(\nu)<\infty.$
Moreover, $D_{r}(\nu)=\xi_{r}$ increases to the box-counting dimension of $F$ as $r\to\infty$, and decreases to the Hausdorff dimension of $\nu$ as $r\to 0$ (cf. [7]).
For every $n\geq 1$, let $\mathcal{C}_{n}$ denote the partition of $\mathbb{R}^{d}$ by cubes of the form $\prod_{h=1}^{d}[k_{h}2^{-n},(k_{h}+1)2^{-n})$ with $(k_{h})_{h=1}^{d}\in\mathbb{Z}^{d}$. For a Borel measure $\nu$, we denote its topological support by $K_{\nu}$. We define
$\mathcal{C}_{n}^{\flat}(\nu):=\{C\in\mathcal{C}_{n}:C\cap K_{\nu}\neq\emptyset\}.$
For every $q\geq 0$, the $L^{q}$-spectrum $T_{\nu}(q)$ for $\nu$ can be defined by (cf. [4, 23])
$T_{\nu}(q):=\lim_{n\to\infty}\frac{\log\sum_{C\in\mathcal{C}_{n}^{\flat}(\nu)}\nu(C)^{q}}{n\log 2},$
if the limit exists; otherwise, one can consider the upper and lower limits. In the following, we simply write $T(q)$ for $T_{\nu}(q)$, because no confusion could arise. For $q\neq 1$, the Rényi dimension for $\nu$ is defined by $R_{\nu}(q):=T(q)/(1-q)$.
Kesseböhmer et al identified the upper quantization dimension for an arbitrary compactly supported measure with its Rényi dimension at a critical point and provided several sufficient conditions that guarantee the existence of the quantization dimension (cf. [14, Theorem 1.1]). This work, along with Feng-Wang’s results in [4], yields that the quantization dimensions exist for the self-affine measures on a large class of planar self-affine sets, including Lalley-Gatzouras carpets. However, the results in [14] do not provide us with useful information on the asymptotic order for the quantization error. In the present paper, we will determine the exact convergence order of the quantization error for the self-affine measures on Lalley-Gatzouras carpets in general.
@@ 1.3. Lalley-Gatzouras carpets
Let $G:=\{(i,j):\;1\leq i\leq n_{j},1\leq j\leq m\}$ and $G_{y}:=\{1,\ldots,m\}$. We consider the following mappings on $\mathbb{R}^{2}$:
(1.3) $f_{ij}(x,y)=(x,y)\left(\begin{array}[]{cccc}a_{ij}&0\\ 0&b_{j}\\ \end{array}\right)+(c_{ij},d_{j}),\;(i,j)\in G.$
With the assumptions (A1)-(A4), the iterated function system $\{f_{ij}\}_{(i,j)\in G}$ satisfies the OSC with respect to $U:=(0,1)^{2}$. By [11], there exists a unique non-empty compact set $E$ satisfying
@@
We call $\mu$ the self-affine measure associated with $\mathcal{I}$ and $\mathcal{P}$. Let $E_{0}:=[0,1]^{2}$. For $l\geq 1$ and $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\in G^{l}$, we write $f_{\sigma}:=f_{i_{1}j_{1}}\circ\cdots\circ f_{i_{l}j_{l}}$. We call $f_{\sigma}(E_{0})$ a cylinder of order $l$.
In the past decades, self-affine sets and self-affine measures have attracted great interest of mathematicians (cf. [1, 4, 13, 15, 16, 17, 19, 20, 21, 22]). As was noted in the literature, problems concerning typical self-affine sets and self-affine measures are usually difficult. In [4], D.-J. Feng and Y. Wang determined various dimensions for a class of planar self-affine sets and self-affine measures, and established several computable formulas for such measures. We refer to [16, 21] for related results on self-affine measures on sponges in $\mathbb{R}^{d}$.
@@
$\widetilde{G}\subset\{0,1,\ldots,n_{0}-1\}\times\{0,1,\ldots,m_{0}-1\};$
$a_{ij}\equiv\frac{1}{n_{0}}\;((i,j)\in\widetilde{G}),\;b_{j}\equiv\frac{1}{m_{0}}\;(j\in\widetilde{G}_{y}),$
$c_{ij}\equiv\frac{i}{n_{0}}\;((i,j)\in\widetilde{G}),\;d_{j}\equiv\frac{j}{m_{0}}\;(j\in\widetilde{G}_{y});$
where $\widetilde{G}_{y}$ denotes the projection of $\widetilde{G}$ onto the $y$-axis. Then the carpet $E$ degenerates to a Bedford-McMullen carpet. Let $\mu$ be as defined in (1.4) with $\widetilde{G}$ in place of $G$. Let $\xi=[\frac{\log m_{0}}{\log n_{0}}]$ and $q_{j}:=\sum_{i:(i,j)\in\widetilde{G}}p_{ij}$. Kesseböhmer and Zhu proved that $D_{r}(\mu)=d_{r}$, where $d_{r}$ is given by
(1.5) $\bigg(\sum_{(i,j)\in\widetilde{G}}(p_{ij}m_{0}^{-r})^{\frac{d_{r}}{d_{r}+r}}\bigg)^{\xi}\bigg(\sum_{j\in\widetilde{G}_{y}}(q_{j}m_{0}^{-r})^{\frac{d_{r}}{d_{r}+r}}\bigg)^{1-\xi}=1.$
@@
Thanks to the relatively fine structure of Bedford-McMullen carpets, we considered the auxiliary coding space $\widetilde{\Phi}_{\infty}:=\widetilde{G}^{\mathbb{N}}\times\widetilde{G}_{y}^{\mathbb{N}}$ and constructed a Bernoulli product $W$ as an auxiliary measure. Then the upper and lower quantization coefficient for $\mu$ were well estimated via the measure $W$, by going back and forth between $E$ and $\Phi_{\infty}$. In the Lalley-Gatzouras case, however, neither the linear parts nor the translations of the mappings $f_{ij},(i,j)\in G$, need to be constant. This substantially adds to the complexity of the structure of the carpets and the difficulty in analyzing the quantization error.
Combining [4, Theorem 2] and [13, Theorem 1.11], we obtain that, the quantization dimension exists for an arbitrary self-affine measure on Lalley-Gatzouras carpets; and the quantization dimension can be expressed in terms of the $L^{q}$-spectrum of the projection of $\mu$ onto the $y$-axis. However, such expressions seem to be inconvenient in the study of the asymptotic order of the quantization error for $\mu$. We will establish a new expression for the quantization dimension in terms of cylinders of the same order. This expression is more convenient for us to construct suitable auxiliary measures which will be used to estimate the quantization coefficient.
@@ 1.4. Some notations
For $\sigma=((i_{1},j_{1}),\ldots,(i_{k},j_{k}))\in G^{k}$ and $1\leq h\leq k$, we define
$|\sigma|=k,\;\sigma|_{h}:=((i_{1},j_{1}),\ldots,(i_{h},j_{h})).$
Let $k\geq 1$. Let $\theta$ denote the empty word. For $\sigma\in G^{1}$, we define $\sigma^{-}=\theta$. For every $k\geq 2$ and $\sigma\in G^{k}$, we define $\sigma^{-}:=\sigma|_{k-1}$. Let $G^{*}:=\bigcup_{k\geq 1}G^{k}$. If $\sigma,\omega\in G^{*}\cup G^{\mathbb{N}}$ satisfy $\sigma=\omega|_{|\sigma|}$, we say that $\sigma$ is comparable with $\omega$ and write $\sigma\preceq\omega$. If $\sigma\preceq\omega$ and $\sigma\neq\omega$, we write $\sigma\precneqq\omega$. We call $\sigma,\omega\in G^{*}$ incomparable if neither $\sigma\preceq\omega$ nor $\omega\preceq\sigma$. For the words $\tau$ in $G_{y}^{*}:=\bigcup_{k\geq 1}G_{y}^{k}$, we define $\tau|_{h},\tau^{-}$ and the partial order $\preceq$, in a similar manner.
@@
$\omega_{y}:=(j_{1},\ldots,j_{l}),\;(\omega\ast\tau)_{y}=(j_{1},\ldots,j_{k});$
$a_{\omega}:=\prod_{h=1}^{l}a_{i_{h}j_{h}},\;b_{\tau}=\prod_{h=l+1}^{k}b_{j_{h}};\;a_{\theta}=b_{\theta}:=1.$
@@
$B_{L,\sigma}:=d_{j_{1}}+\sum_{p=2}^{k}\bigg(\prod_{h=1}^{p-1}b_{j_{h}}\bigg)d_{j_{p}},\;B_{R,\sigma}:=B_{L,\sigma}+\prod_{h=1}^{k}b_{j_{h}}.$
@@
For $\sigma,\omega\in\Psi^{*}$, we write $\sigma\preceq\omega$ if $F_{\sigma}\supset F_{\omega}$. In Lalley-Gatzouras case, the diameters of approximate squares of the same order can be widely different.
One can see that $\Psi^{*}$ provides location codes for approximate squares. However, the way how an approximate square $F_{\sigma}$ has its sub-approximate squares heavily depends on $\sigma$. This leads to great difficulty in constructing auxiliary measures on the carpets $E$. Therefore, we will also consider the auxiliary coding space $\Phi_{\infty}=G^{\mathbb{N}}\times G_{y}^{\mathbb{N}}$. Although $\Phi_{\infty}$ is not completely compatible with the carpet $E$ (cf. Remark 3.3), it suits our purpose perfectly. We define
@@
We write $\mathcal{L}(\sigma)_{L}:=\sigma_{L}$ and $\mathcal{L}(\sigma)_{R}:=\sigma_{R}$. For $\sigma,\omega\in\Phi^{*}$, we write $\sigma\preceq\omega$, if $\sigma_{L}\preceq\omega_{L}$ and $\sigma_{R}\preceq\omega_{R}$. We call $\sigma,\omega\in\Phi^{*}$ incomparable if neither $\sigma\preceq\omega$ nor $\omega\preceq\sigma$. For every $\sigma\in\Phi^{*}$, we define $\sigma]:=[\sigma_{L}]\times[\sigma_{R}]$, where
@@
The main difference between $\Phi^{*}$ and $\Phi^{*}$ lies in the fact that they are endowed with different partial orders. Some more information will be presented in Remark 3.3.
@@
Let $|B|$ denote the diameter of a set $B\subset\mathbb{R}^{2}$ and $B^{\circ}$ its interior. We have the following facts.
@@
    For $\sigma,\omega\in\Psi^{*}$, we have, either $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset$, or $F_{\sigma}\subset F_{\omega}$, or $F_{\omega}\subset F_{\sigma}$.